how do you solve this using substitution or elimination:
2x=-7+3y
3x-4y=-9

Answer by ewatrrr(24785)


Hi,
*Note: either substitution or elimination would work. Sets up well for elimination
2x=-7+3y
2x - 3y = -7
2x - 3y = -7
3x - 4y = -9
-8x + 12y = 28 Mulitplying both sides of the first equation by -4
9x - 12y = -27 Mulitplying both sides of the second equation by 3
x = 1
2*1 = -7+ 3y
9 = 3y
y = 3